# Conceptualization, Methodology, Writing – original draft, Software, Writing – review & editing

### v1

@@ Abstract
Real-time prediction of corridor-level crash occurrence risk (COR) remains challenging, as existing near-miss based extreme value models oversimplify collision geometry, exclude vehicle–infrastructure (V–I) interactions, and inadequately capture spatial heterogeneity in vehicle dynamics. This study introduces a geometry-aware two-dimensional time-to-collision (2D-TTC) indicator within a Hierarchical Bayesian spatiotemporal grouped random parameter (HBSGRP) framework using a non-stationary univariate generalized extreme value (UGEV) model to estimate short-term COR in urban corridors. High-resolution trajectories from the Argoverse-2 dataset, covering 28 locations along Miami’s Biscayne Boulevard, were analyzed to extract extreme V–V and V–I near misses. The model incorporates dynamic variables and roadway features as covariates, with partial pooling across locations to address unobserved heterogeneity. Results show that the HBSGRP–UGEV framework outperforms fixed-parameter alternatives, reducing DIC by up to 7.5% for V–V and 3.1% for V–I near-misses. Predictive validation using ROC–AUC confirms strong performance: 0.89 for V–V segments, 0.82 for V–V intersections, 0.79 for V–I segments, and 0.75 for V–I intersections. Model interpretation reveals that relative speed and distance dominate V–V risks at intersections and segments, with deceleration critical in segments, while V–I risks are driven by speed, boundary proximity, and steering/heading adjustments. These findings highlight the value of a statistically rigorous, geometry-sensitive, and spatially adaptive modeling approach for proactive corridor-level safety management, supporting real-time interventions and long-term design strategies aligned with Vision Zero.
@@ 1 Introduction
The increasing complexity of urban road networks, rising traffic volumes, and the coexistence of human-driven vehicles (HDVs) and autonomous vehicles (AVs) demand innovative safety assessment approaches. Urban corridors are particularly vulnerable, where a single incident can disrupt mobility, trigger cascading congestion, and impose severe economic losses. In 2023, traffic crashes in the United States claimed 40,901 lives, with a fatality rate of 1.26 per 100 million vehicle miles traveled (VMT) (National Center for Statistics and Analysis, 2025). The total annual cost of crashes is estimated to exceed \$1.85 trillion, including \$460 billion in direct economic losses and \$1.4 trillion in reduced quality of life (Blincoe et al., 2023). Despite extensive research on crash patterns, risk estimation, and mitigation strategies (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b), accurate prediction of crash risk in urban corridors remains methodologically underdeveloped.
@@
Most existing models focus exclusively on vehicle–vehicle (V–V) near-misses, largely due to the availability of vehicle trajectory data, while systematically excluding vehicle–infrastructure (V–I) interactions such as lane edges, curbs, barriers, and crosswalks. Yet V–I events account for a substantial proportion of severe run-off-road and roadway departure crashes (Islam et al., 2021). This omission leaves corridor-level safety models inherently incomplete. Recent advances in AV sensing and high-definition (HD) mapping, such as the Argoverse-2 dataset (Wilson et al., 2023), now enable the precise integration of roadway boundaries with vehicle trajectories, allowing both V–V and V–I near-misses to be modeled with geometric fidelity.
@@
Crash risk prediction strategies based on historical retrospective crash data (Arun et al., 2021; Lee et al., 2017; Pei et al., 2011) suffer from well-documented limitations (Lord and Mannering, 2010; Lord et al., 2021; Mannering et al., 2020), including the rarity and stochastic nature of crashes, underreporting (Arun et al., 2021; Tarko, 2018), temporal delays, and spatial mismatches with real-time conditions. These challenges motivate a paradigm shift toward proactive, trajectory-based surrogate safety measures (SSMs) (Ali et al., 2023a; Arun et al., 2021; Mahmud et al., 2017). Derived from high-resolution vehicle trajectory data, SSMs reflect underlying near-miss dynamics (Li et al., 2024; Tarko, 2018; Davis et al., 2011). Among SSMs, time-based indicators such as Time-to-Collision (TTC), Post-Encroachment Time (PET), and Modified TTC (MTTC) are widely used (Hayward, 1971; Perkins and Harris, (1967; Minderhoud and Bovy, 2001; Ozbay et al., 2008; Allen et al., 1978; Venthuruthiyil and Chunchu, 2022). However, these formulations often assume constant velocity and linear trajectories, which bias results in boundary-constrained environments. Recent research has extended TTC into higher-fidelity two-dimensional (2D) indicators (Li et al., 2024; Anis et al., 2025b), incorporating latitude and longitude position, heading, and acceleration. Yet, even these advances typically employ simplified circular approximations, which can influence risk estimates in urban contexts where vehicle shape and alignment significantly influence near-miss likelihood and severity. To support reliable short-term crash occurrence risk (COR) estimation, SSM frameworks must enhance geometric fidelity to more accurately represent real-world interactions between vehicles and infrastructure. This study builds on these efforts by introducing a high-fidelity, geometry-aware 2D-TTC indicator that explicitly incorporates vehicle geometry, dynamic states, and roadway boundaries, ensuring realistic and efficient detection of both V–V and V–I near misses.
@@
Early EVT applications in traffic safety employed stationary GEV models (Wang et al., 2018; Guo et al., 2019; Orsini et al., 2019; Alozi and Hussein, 2022; Hussain et al., 2022) that assumed constant parameterization across the study area. More recently, non-stationary GEV models have been introduced, allowing covariates to influence location and scale parameters (Tahir and Haque, 2024; Ali et al., 2022; Fu and Sayed, 2022a; Zheng et al., 2019; Nazir et al., 2023; Kar et al., 2024; Ali et al., 2023b; Alozi and Hussein, 2022; Zheng et al., 2014; Kar et al., 2023) and thereby capture heterogeneity across traffic states (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b). Bayesian hierarchical extensions have further addressed data sparsity and unobserved heterogeneity (Zheng and Sayed, 2019; Fu and Sayed, 2022b). However, most applications remain geographically narrow and functionally limited to either intersections (Fu and Sayed, 2023, 2021c, 2022b) or segments (Kumar and Mudgal, 2024; Kamel et al., 2023; Kamel and Sayed, 2024; Anis et al., 2025b; Kamel et al., 2024). Although some studies have incorporated vehicle dynamics (Kumar and Mudgal, 2024; Fu and Sayed, 2022b; Anis et al., 2025b) and examined heterogeneity across traffic flow directions or roadway geometries (Kamel et al., 2023, 2024), these frameworks generally omit explicit integration of roadway geometry and directional heterogeneity at the corridor level. Such integration requires a modeling structure capable of jointly representing vehicle dynamics (which vary temporally) and roadway features (which vary spatially across segments and intersections). Therefore, a critical gap remains for a comprehensive Bayesian hierarchical GEV framework capable of jointly modeling vehicle dynamics, roadway geometry, flow directionality, and multilevel heterogeneity. Such a framework is essential for accurate corridor-level COR estimation that reflects the complex spatiotemporal dynamics of urban traffic environments. This paper makes the following unique contributions:
  • Develops a corridor-level COR framework that simultaneously incorporates both V–V and V–I interactions, addressing a critical omission in prior studies.
  • Introduces a high-fidelity, geometry-aware 2D-TTC indicator that realistically captures extreme near-miss events by accounting for vehicle geometry, dynamic states, and roadway boundaries.
@@ 2 Near-miss detection framework
This section presents the modeling framework for computing a geometry-aware, high-fidelity 2D-TTC to support COR estimation. The framework integrates a continuous-time kinematic bicycle model, vehicle geometry representation, and a spatial near-miss detection algorithm, allowing capture of V-V and V-I interactions. These components collectively enable the accurate forecasting of vehicle trajectories and the identification of potential near misses. The resulting architecture forms the computational backbone of the SSM analysis, ensuring a realistic representation of near-miss events in mixed-traffic environments.
@@ 2.2 Vehicle geometry representation
While the dynamic model predicts vehicle positions and headings, accurate near-miss detection requires a realistic representation of the vehicle footprint. Simplifying vehicles as points or circles often misrepresents interactions, especially during lane changes, swerving, or angled approaches. To address this, each vehicle is modeled as a rigid rectangle defined by its actual length $L_{i}^{\text{veh}}$ and width $W_{i}^{\text{veh}}$ (vehicle $i\in\{A,B\}$). The vehicle is centered at its center of gravity, with four corners specified in the body-fixed coordinate system as:
@@ 2.3 V-V near-miss detection
Vehicles are modeled as rigid rectangles, and their interactions are determined by predicted states and orientations. At each timestep of the RK4 integration, the global positions of all four corners are computed using the kinematic and rotational transformations. A collision is defined to occur when a corner of one vehicle lies within a proximity threshold of a corner of another. Let ${p}_{A}^{(j)}(t)$ and ${p}_{B}^{(k)}(t)$ denote the global coordinates of corners $j$ and $k$ of vehicles A and B at time $t$. A V–V near-miss is flagged if any pair of these corners satisfies the proximity condition.
@@ 2.4 V-I near-miss detection
In addition to V-V interactions, the framework also detects near-misses between vehicles and roadway infrastructure such as curbs, medians, or barriers. Capturing these V–I events is critical for identifying edge encroachments or departures that often precede run-off-road crashes.
@@ 15         Record V-I near-miss, set $t_{c}$;
While this geometric framework enables precise localization of near-misses, it does not capture the probabilistic structure or temporal evolution of COR. To address this, the extracted blockwise minima of 2D-TTC are modeled using the EVT model. For each block, the model estimates the probability that 2D-TTC falls below a critical safety threshold, which defines the COR metric as a short-term indicator of crash risk. Block duration is chosen to balance temporal sensitivity with statistical stability across diverse contexts. This framework, therefore, combines deterministic near-miss detection with probabilistic EVT modeling to enable localized, data-driven assessment of crash-prone conditions. The complete methodological pipeline, from trajectory simulation to dynamic risk estimation, is summarized in Fig. 2.5.
@@ 3 Extreme value theory (EVT)
Extreme Value Theory (EVT) provides the statistical foundation for estimating COR by extrapolating rare, high-severity events from frequent near-misses. In this study, extreme events are represented by block maxima of 2D-TTC derived from high-resolution vehicle trajectories (Algorithm 1). EVT allows risk estimation without requiring large volumes of historical crash data (Zheng and Sayed, 2020; Fu and Sayed, 2022a). Two EVT approaches are commonly applied: the Block Maxima (BM) method, which fits maxima or minima from fixed-duration blocks using the Generalized Extreme Value (GEV) distribution (Coles et al., 2001), and the Peaks Over Threshold (POT) method, which models threshold exceedances using the Generalized Pareto Distribution (GPD). While POT can be data-efficient, it is highly sensitive to threshold choice and less compatible with discrete, sensor-derived datasets. By contrast, BM aligns naturally with real-time AV data, avoids subjective thresholds, and supports scalable, time-series analysis (Fu and Sayed, 2022a; Songchitruksa and Tarko, 2006). Accordingly, this study adopts the BM–GEV framework (see Fig. 2.5).
The GEV has been widely applied in traffic safety research. Early studies modeled near-missets with univariate and multivariate GEV formulations (Orsini et al., 2020; Zhang et al., 2019; Zheng et al., 2014). Recent advances incorporate nonstationary structures, linking GEV parameters to covariates such as vehicle dynamics (Anis et al., 2025b; Fu and Sayed, 2021c, 2022b). Bayesian hierarchical models have further addressed sparse data and unobserved heterogeneity across vehicles, sites, and traffic contexts. The increasing availability of high-resolution AV trajectory and HD map data has highlighted the need for flexible GEV formulations that jointly capture V–V and V–I interactions.
Despite this progress, most UGEV models still neglect the joint influence of roadway geometry and vehicle dynamics, two critical but distinct sources of heterogeneity. Roadway geometry (e.g., lane count, curvature, median type) is spatially structured and relatively stable within segments or intersections, while vehicle dynamics (e.g., speed, acceleration) vary temporally and spatially, driving short-term behavioral volatility. Capturing this multilevel heterogeneity is crucial for estimating corridor- and network-level crash risk. To address this gap, this study develops a Hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) UGEV model. Unlike standard random-parameter models, which assign variation independently to each unit, grouped random parameters capture shared variation across clusters (e.g., segment types, corridor directions) while allowing for within-group variability. This structure provides both flexibility and parsimony, making it well-suited for high-resolution, corridor-scale crash risk modeling (Washington et al., 2020; Cai et al., 2018; Islam et al., 2023).

### v2

@@ Abstract
Real-time prediction of corridor-level crash occurrence risk (COR) remains challenging, as existing near-miss based extreme value models oversimplify collision geometry, exclude vehicle–infrastructure (V–I) interactions, and inadequately capture spatial heterogeneity in vehicle dynamics. This study introduces a geometry-aware two-dimensional time-to-collision (2D-TTC) indicator within a Hierarchical Bayesian spatiotemporal grouped random parameter (HBSGRP) framework using a nonstationary univariate generalized extreme value (UGEV) model to estimate short-term COR in urban corridors. High-resolution trajectories from the Argoverse-2 dataset, covering 28 locations along Miami’s Biscayne Boulevard, were analyzed to extract extreme near misses (V–V and V–I). The model incorporates dynamic variables and road features as covariates, with partial pooling across locations to address unobserved heterogeneity. The results show that the HBSGRP–UGEV framework outperforms fixed-parameter alternatives, reducing the DIC by up to 7.5% for V–V and 3.1% for V–I near-misses. Predictive validation using ROC–AUC confirms strong performance: 0.89 for V–V segments, 0.82 for V–V intersections, 0.79 for V–I segments, and 0.75 for V–I intersections. Model interpretation reveals that relative speed and distance dominate V–V risks at intersections and segments, with deceleration critical in segments, while V–I risks are driven by speed, boundary proximity, and steering/heading adjustments. These findings highlight the value of a statistically rigorous, geometry-sensitive, and spatially adaptive modeling approach for proactive corridor-level safety management, supporting real-time interventions and long-term design strategies aligned with Vision Zero.
@@ 1 Introduction
The increasing complexity of urban road networks, rising traffic volumes, and the coexistence of human-driven vehicles (HDVs) and autonomous vehicles (AVs) require innovative safety assessment approaches. Urban corridors are particularly vulnerable, where a single incident can disrupt mobility, trigger cascading congestion, and impose severe economic losses. In 2023, traffic crashes in the United States claimed 40901 lives, with a fatality rate of 1.26 per 100 million vehicle miles traveled (VMT) (National Center for Statistics and Analysis, 2025). The total annual cost of crashes is estimated to exceed \$1.85 trillion, including \$460 billion in direct economic losses and \$1.4 trillion in reduced quality of life (Blincoe et al., 2023). Despite extensive research on crash patterns, risk estimation, and mitigation strategies (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b), accurate prediction of crash risk in urban corridors remains methodologically underdeveloped.
@@
Most existing models focus exclusively on vehicle–vehicle (V–V) near-misses, largely due to the availability of vehicle trajectory data, while systematically excluding vehicle–infrastructure (V–I) interactions such as lane edges, curbs, barriers, and crosswalks. However, V–I events account for a substantial proportion of severe run-off-road and roadway departure crashes (Islam et al., 2021). This omission makes corridor-level safety models inherently incomplete. Recent advances in AV sensing and high-definition (HD) mapping, such as the Argoverse-2 dataset (Wilson et al., 2023), now allow the precise integration of roadway boundaries with vehicle trajectories, allowing both V–V and V–I near-misses to be modeled with geometric fidelity.
@@
Crash risk prediction strategies based on historical retrospective crash data (Arun et al., 2021; Lee et al., 2017; Pei et al., 2011) suffer from well-documented limitations (Lord and Mannering, 2010; Lord et al., 2021; Mannering et al., 2020), including the rarity and stochastic nature of crashes, underreporting (Arun et al., 2021; Tarko, 2018), temporal delays, and spatial mismatches with real-time conditions. These challenges motivate a paradigm shift towards proactive trajectory-based surrogate safety measures (SSMs) (Ali et al., 2023a; Arun et al., 2021; Mahmud et al., 2017). Using high-resolution vehicle trajectory data, SSMs reflect the underlying near-miss dynamics (Li et al., 2024; Tarko, 2018; Davis et al., 2011). Among SSMs, time-based indicators such as Time-to-Collision (TTC), Post-Encroachment Time (PET), and Modified TTC (MTTC) are widely used (Hayward, 1971; Perkins and Harris, (1967; Minderhoud and Bovy, 2001; Ozbay et al., 2008; Allen et al., 1978; Venthuruthiyil and Chunchu, 2022). However, these formulations often assume constant velocity and linear trajectories, leading to biased results in boundary-constrained environments. Recent research has extended TTC to two-dimensional (2D) indicators of higher-fidelity (Li et al., 2024; Anis et al., 2025b), incorporating latitude and longitude position, heading, and acceleration. However, even these advances typically employ simplified circular approximations, which can influence risk estimates in urban contexts where vehicle shape and alignment significantly influence the likelihood and severity of near misses. To support a reliable short-term crash occurrence risk (COR) estimate, SSM frameworks must enhance geometric fidelity to more accurately represent real-world interactions between vehicles and infrastructure. This study builds on these efforts by introducing a high-fidelity, geometry-aware 2D-TTC indicator that explicitly incorporates vehicle geometry, dynamic states, and roadway boundaries, ensuring realistic and efficient detection of both V–V and V–I near misses.
@@
Early applications of EVT in traffic safety employed stationary GEV models (Wang et al., 2018; Guo et al., 2019; Orsini et al., 2019; Alozi and Hussein, 2022; Hussain et al., 2022) that assumed constant parameterization across the study area. More recently, non-stationary GEV models have been introduced, allowing covariates to influence location and scale parameters (Tahir and Haque, 2024; Ali et al., 2022; Fu and Sayed, 2022a; Zheng et al., 2019; Nazir et al., 2023; Kar et al., 2024; Ali et al., 2023b; Alozi and Hussein, 2022; Zheng et al., 2014; Kar et al., 2023) and thereby capture heterogeneity across traffic states (Lanzaro et al., 2023; Kamel et al., 2023, 2024; Kamel and Sayed, 2024; Singh et al., 2024; Ghoul and Sayed, 2025; Singh et al., 2025; Anis et al., 2025b). Bayesian hierarchical extensions have also addressed data sparsity and unobserved heterogeneity (Zheng and Sayed, 2019; Fu and Sayed, 2022b). However, most applications remain geographically narrow and functionally limited to either intersections (Fu and Sayed, 2023, 2021c, 2022b) or segments (Kumar and Mudgal, 2024; Kamel et al., 2023; Kamel and Sayed, 2024; Anis et al., 2025b; Kamel et al., 2024). Although some studies have incorporated vehicle dynamics (Kumar and Mudgal, 2024; Fu and Sayed, 2022b; Anis et al., 2025b) and examined heterogeneity across traffic flow directions or roadway geometries (Kamel et al., 2023, 2024), these frameworks generally omit explicit integration of roadway geometry and directional heterogeneity at the corridor level. Such integration requires a modeling structure capable of jointly representing vehicle dynamics (which vary temporally) and roadway features (which vary spatially across segments and intersections). Therefore, a critical gap remains for a comprehensive Bayesian hierarchical GEV framework capable of jointly modeling vehicle dynamics, roadway geometry, flow directionality, and multilevel heterogeneity. This framework is crucial for accurately estimating the COR at the corridor level, which reflects the complex spatiotemporal dynamics of urban traffic environments. This paper makes the following unique contributions:
  • Develops a corridor-level COR framework that simultaneously incorporates both V–V and V–I interactions, addressing a critical omission in previous studies.
  • Introduces a geometry-aware high-fidelity 2D-TTC indicator that realistically captures extreme near-miss events by accounting for vehicle geometry, dynamic states, and roadway boundaries.
@@ 2 Near-miss detection framework
This section presents the modeling framework for computing a geometry-aware, high-fidelity 2D-TTC to support COR estimation. The framework integrates a continuous-time kinematic bicycle model, vehicle geometry representation, and a spatial near-miss detection algorithm, allowing capture of V-V and V-I interactions. Together, these components enable the accurate forecasting of vehicle trajectories and identification of potential near misses. The resulting architecture forms the computational backbone of the SSM analysis, ensuring a realistic representation of near-miss events in mixed-traffic environments.
@@ 2.2 Vehicle geometry representation
Although the dynamic model predicts vehicle positions and headings, accurate near-miss detection requires a realistic representation of the vehicle footprint. Simplifying vehicles as points or circles often misrepresents interactions, especially during lane changes, swerving, or angled approaches. To address this, each vehicle is modeled as a rigid rectangle defined by its actual length $L_{i}^{\text{veh}}$ and width $W_{i}^{\text{veh}}$ (vehicle $i\in\{A,B\}$). The vehicle is centered at its center of gravity, with four corners specified in the body-fixed coordinate system as:
@@ 2.3 V-V near-miss detection
Vehicles are modeled as rigid rectangles, and their interactions are determined by predicted states and orientations. At each timestep of the RK4 integration, the global positions of all four corners are computed using the kinematic and rotational transformations. A collision is defined when a vehicle’s corner lies within a proximity threshold of a corner of another vehicle. Let ${p}_{A}^{(j)}(t)$ and ${p}_{B}^{(k)}(t)$ denote the global coordinates of corners $j$ and $k$ of vehicles A and B at time $t$. A V–V near-miss is flagged if any pair of these corners satisfies the proximity condition.
@@ 2.4 V-I near-miss detection
In addition to V-V interactions, the framework also detects near misses between vehicles and roadway infrastructure, including curbs, medians, or barriers. Capturing these V–I events is critical for identifying edge encroachments or departures that often precede run-off-road crashes.
@@ 15         Record V-I near-miss, set $t_{c}$;
While this geometric framework enables precise localization of near misses, it does not capture the probabilistic structure or temporal evolution of COR. To address this, the extracted blockwise minima of 2D-TTC are modeled using the EVT model. For each block, the model estimates the probability that 2D-TTC falls below a critical safety threshold, which defines the COR metric as a short-term indicator of crash risk. Block duration is chosen to balance temporal sensitivity with statistical stability across diverse contexts. This framework, therefore, combines deterministic near-miss detection with probabilistic EVT modeling to enable localized, data-driven assessment of crash-prone conditions. The complete methodological pipeline, from trajectory simulation to dynamic risk estimation, is summarized in Fig. 2.5.
@@ 3 Extreme value theory (EVT)
Extreme Value Theory (EVT) provides the statistical foundation for estimating COR by extrapolating rare, high-severity events from frequent near-misses. In this study, extreme events are represented by block maxima of 2D-TTC derived from high-resolution vehicle trajectories (Algorithm 1). EVT allows risk estimation without requiring large volumes of historical crash data (Zheng and Sayed, 2020; Fu and Sayed, 2022a). Two EVT approaches are commonly applied: the Block Maxima (BM) method, which fits maxima or minima from fixed-duration blocks using the Generalized Extreme Value (GEV) distribution (Coles et al., 2001), and the Peaks Over Threshold (POT) method, which models threshold exceedances using the Generalized Pareto Distribution (GPD). Although POT can be data-efficient, it is highly sensitive to threshold choice and less compatible with discrete sensor-derived datasets. By contrast, BM aligns naturally with real-time AV data, avoids subjective thresholds, and supports scalable, time-series analysis (Fu and Sayed, 2022a; Songchitruksa and Tarko, 2006). Consequently, this study adopts the BM–GEV framework (see Fig. 2.5).
The GEV has been widely applied in traffic safety research. Early studies modeled near-missets with univariate and multivariate GEV formulations (Orsini et al., 2020; Zhang et al., 2019; Zheng et al., 2014). Recent advances incorporate nonstationary structures, linking GEV parameters to covariates such as vehicle dynamics (Anis et al., 2025b; Fu and Sayed, 2021c, 2022b). Bayesian hierarchical models have also addressed sparse data and unobserved heterogeneity across vehicles, sites, and traffic contexts. The increasing availability of high-resolution AV trajectory and HD map data has highlighted the need for flexible GEV formulations that jointly capture V–V and V–I interactions.
Despite this progress, most UGEV models still neglect the joint influence of roadway geometry and vehicle dynamics, two critical but distinct sources of heterogeneity. Roadway geometry (e.g., lane count, curvature, median type) is spatially structured and relatively stable within segments or intersections, while vehicle dynamics (e.g., speed, acceleration) vary temporally and spatially, driving short-term behavioral volatility. Capturing this multilevel heterogeneity is crucial for estimating corridor and network level crash risk. To address this gap, this study develops a hierarchical Bayesian Spatiotemporal Grouped Random Parameter (HBSGRP) UGEV model. Unlike standard random-parameter models, which assign variation independently to each unit, grouped random parameters capture shared variation across clusters (e.g., segment types, corridor directions) while allowing for within-group variability. This structure provides both flexibility and parsimony, making it well-suited for high-resolution, corridor-scale crash risk modeling (Washington et al., 2020; Cai et al., 2018; Islam et al., 2023).